### v2

# Using Machine Learning to Explore Defect Configurations in Cd/Zn-Se/Te Compounds

Abstract
We developed a framework for predicting the energies and ground state configurations of native point defects, extrinsic dopants and impurities, and defect complexes across zincblende-phase Cd/Zn-Te/Se/S compounds, important for CdTe-based solar cells. This framework, named DeFecT-FF, is powered by high-throughput density functional theory (DFT) computations and crystal graph-based machine learning force field (MLFF) models trained on the DFT data. The Cd/Zn-Te/Se/S chemical space is chosen because alloying at Cd or Te sites is a promising avenue to tailor the electronic and defect properties of the CdTe absorber layer to potentially improve solar cell performance. The sheer number of defect configurations achievable when considering all possible singular defects and their combinations, symmetry-breaking operations, and defect charge states, as well as the expense of running large supercell calculations, makes this an ideal problem for developing accurate and widely-applicable force field models. Here, we introduce our dataset of structures and energies from HSE06 geometry optimization, including bulk and alloyed supercells with and without defects. Data were gradually expanded using active learning and accurate MLFF models were trained to predict energies and atomic forces across different charge states. Via accelerated prediction and screening, we identified many new low energy defect configurations and obtained high-fidelity defect formation energy diagrams using HSE06 calculations with spin-orbit coupling. The DeFecT-FF framework has been released publicly as an online tool on the nanoHUB platform, allowing users to upload any crystallographic information file, generate defects of interest, and compute defect formation energies as a function of Fermi level and chemical potential conditions, thus bypassing expensive DFT calculations.
DeFecT‑FF: Accelerated Modeling of Defects in Cd/Zn-Te/Se/S Compounds Combining High-Throughput DFT and Machine Learning Force Fields
Md Habibur Rahmana, Maitreyo Biswasa and Arun Mannodi-Kanakkithodia
footnotetext: aSchool of Materials Engineering, Purdue University, West Lafayette, IN 47907, USA; E-mail: amannodi@purdue.edu
1 Introduction
Advancements in solar cell technologies are crucial for meeting global energy demands and supporting the transition to a decarbonized grid Srisuda1, Srisuda2, a2bcx4_abx2, 11133092. Among photovoltaic (PV) technologies, CdTe ranks second after crystalline Si, accounting for about 7% of the global market CdTe_1, Srisuda1, Srisuda2. Its commercial success arises from a direct bandgap of $\sim$1.5 eV, high absorption coefficient ($5\times 10^{5}$ cm-1) CdTe_2, eduardo_men_ndez_proupin_dae1679b, fadhil_k__alfadhili_1e77f8db, o__de_melo_1bf7fe05, kaiying_luo_60e9a185, wei_chao_chen_652bd27c, kai_shen_86d534d1, jinshui_miao_634f66d0, a_gayt_n_garc_a_23ea61bf, xin_zheng_02d025a7, low production cost, and good thin-film conductivity CdTe_3. However, its maximum efficiency of 22.3% remains below the $\sim$30% theoretical limit, mainly due to Shockley-Read-Hall (SRH) recombination associated with grain boundaries, point defects, and dislocations Freysoldt2014-en. Native defects and impurities create trap states within the bandgap that act as nonradiative recombination centers Ganose_Scanlon_2022, Mannodi-Kanakkithodi_2023, such as Cd vacancies (VCd), which accelerate carrier recombination and can reduce power conversion efficiency by nearly 5% sean_1, Turiansky, aneta_wardak_ab3492d5, p__d__hatton_26561f92, michael_a__scarpulla_043ff246.
CdTe often suffers from low hole density, limiting its PV efficiency Srisuda1, Srisuda2. Cu is commonly introduced as an acceptor dopant via high-temperature CdCl2 annealing, where Cu and Cl diffuse at 1017–1019 cm-3, altering electronic properties Yang_Yin, Srisuda1, Srisuda2. While Cui and ClTe act as shallow donors and CuCd as a non-shallow acceptor forming complexes such as (Cui+CuCd) and (Cli+CuCd)2+ Krasikov_1, Krasikov_2, CdTe_1, Cu doping typically yields suboptimal hole density ($\sim$1014 cm-3) compared to the ideal value of $\sim$1016 cm-3 CdTe_3. In contrast, group V dopants such as As achieve higher hole densities without reducing carrier lifetime Ablekim, CdTe_1. Se alloying to create CdSexTe1-x further enhances PV efficiency by improving absorption, band alignment, and carrier lifetimes Ablekim, Fiducia2019-wa. ZnTe, with favorable band alignment, serves as an efficient hole transport layer Gorai_Krasikov. Thus, exploring the defect chemistry of Cd/Zn–S/Se/Te alloys is vital to improving CdTe- and CdSeTe-based thin-film solar cells De_Souza_Harrington_2023.
Defect levels are experimentally measured using cathodoluminescence, photoluminescence, optical spectroscopy, or deep-level transient spectroscopy (DLTS) Lang2003. These methods face significant challenges in sample preparation and assigning measured levels to specific defects Wickramaratne, Kim_Gelczuk. To overcome this, density functional theory (DFT) is widely used to calculate defect formation energy (Ef) as a function of Fermi level (EF), defect charge state (q), and chemical potential ($\mu$) Freysoldt2014-en, Broberg_Bystrom, gw_def, Lee_Din, Grill2004-pk, Buckeridge_2019. DFT enables identification of donor- and acceptor-type defects, shallow or deep defect levels, type of equilibrium conductivity, defect concentrations, and carrier capture rates Mannodi-Kanakkithodi2022-ck, Turiansky, Rahman2023, Defect_Energetics_JPCC, Kim_Park_Hood_Walsh_2019, Lyons2017-sp. When an appropriate level of theory is applied, DFT-computed charge transition levels compare well with experiments Mannodi-Kanakkithodi2022-ck, Kim_Gelczuk, snb_1, snb_2. However, DFT is computationally expensive and scales poorly with system size, making it difficult to explore the vast configurational space of vacancies, interstitials, antisites, and defect complexes across many compounds and charge states mba, Mannodi-HP1.
The prediction of defect properties can be accelerated by integrating DFT simulations with machine learning (ML) approaches such as crystal graph neural networks (GNNs) CGCNN, Choudhary, Chen_Li_Bruna_2017. GNNs effectively represent and predict the energies and properties of molecules, polymers, and crystalline materials Kipf_Welling_2016, Witman_2023 by transforming atomic structures into graphs where atoms are nodes and bonds are edges m3gnet. They learn intricate structural representations to predict properties such as formation or decomposition energy, bandgap, and defect formation energy while reducing computational cost. In prior work, we used GNNs to predict and screen native defects and functional impurities in group IV, III–V, and II–VI zincblende semiconductors Rahman_Gollapalli, covering vacancies, interstitials, anti-site, and extrinsic defects. While the models predicted charge-dependent defect formation energies for different chemical potential conditions, several limitations were observed: (1) training on a broad chemical space (34 compounds) led to large errors for specific compositions; (2) models showed good performance on binaries but lower accuracy for alloy systems such as CdSexTe1-x and CdxZn1-xTe Rahman_Rojsatien, Yang_Yin, Krasikov_2, Srisuda2, Defect_Energetics_JPCC, Rahman_Mannodi-Kanakkithodi_2025b, Lany_soc; (3) reliance on modest 64-atom $2\times 2\times 2$ supercells limited defect complex modeling; (4) use of the semi-local GGA-PBE functional (GGA: Generalized Gradient Approximation, PBE: Perdew–Burke–Ernzerhof) inherently limited prediction fidelity GGA_gaps, HP_hse, Lyons_Van; and (5) dependence on gradient-free optimization with GNN models prevented more efficient gradient-based geometry optimization.
To overcome prior limitations from casting too wide a chemical space, using smaller supercells, and relying on semi-local functionals, we developed a more comprehensive, multi-fidelity methodology. Our approach begins with an initial dataset of bulk and defect configurations spanning Cd/Zn–Te/Se/S binary and multi-nary compounds. These structures were first computed using the PBE functional, providing a baseline set of bulk and defect configurations and their energies, with a substantial portion of the PBE dataset compiled from our previously published works Mannodi-ZB, Mannodi-Kanakkithodi2022-ck. Initial GNN models trained on the PBE dataset  Bapst_Keck served as the foundation for predicting defect properties over a pre-defined defect chemical space containing thousands of vacancies, interstitials, antisites, substitutional defects, and defect complexes. Defect enumeration included all relevant native defects as well as group-V dopants (N, P, As, Sb, Bi) which are promising for achieving p-type conductivity and unintentional impurities such as Cl and O which are known to strongly influence the performance of CdTe and CdSexTe1-x solar cells Li2017-ez, Corsin_2016. The Cd/Zn–Te/Se/S chemical space was chosen due to its relevance to Se grading, Cd–Zn interfaces, and absorber composition tuning in CdTe solar cells, where exploring all low-energy native and extrinsic defects across these compositions provides a comprehensive dataset for experimental comparison. Although ZnS and ZnSe are not primary absorbers, they remain chemically informative, while CdS functions as an important buffer layer CdTe_1, CdTe_2, CdTe_3.
Figure  1: Statistics of the HSE06 dataset: (a) Number of bulk configurations from the CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x, ZnSxSe1-x, and ZnSexTe1-x compositions. (b) Distribution of defect configurations across the Cd–chalcogen and Zn–chalcogen binaries and ternaries. (c) Violin plots of crystal formation energies (meV/atom) for the entire dataset across five charge states (+2+2 to 2-2). (d,e) Defect formation energy diagrams for CdTe under Cd-rich and Te-rich conditions from HSE06 functional, highlighting the relative stability of key native (VCd, VTe) and extrinsic defects (AsTe, ClTe).
To improve prediction accuracy, active learning was employed to iteratively generate new DFT data and refine the GNN models AL_1, AL_2, AL_3, AL_4, AL_5, Gentle_2010, Farache_Verduzco. Following convergence of the active learning scheme, we performed higher-fidelity HSE06 (Heyd–Scuseria–Ernzerhof)  hse calculations on a curated subset of representative PBE-relaxed structures to obtain more accurate bandgaps, charge transition levels, and defect formation energies. GNN models, specifically using the M3GNet m3gnet architecture for machine learning force fields (MLFFs), were then trained on the HSE06 data and subsequently used for new predictions. Our complete methodology is summarized as: PBE data collection $\rightarrow$ initial PBE GNN model training $\rightarrow$ active-learning–driven expansion of the PBE dataset $\rightarrow$ HSE06 refinement of a subset of the PBE dataset $\rightarrow$ training MLFF models at HSE06 accuracy. The next few sections describe our methodology and results in detail, highlighting the following major contributions of this work:
  • Construction of the largest unified HSE06 defect dataset across Cd/Zn–Te/Se/S compositions, including native and extrinsic defects, and defect complexes, simulated in five charge states.
  • Development of an HSE06 MLFF-based defect optimization workflow that is orders of magnitude faster than full DFT.
  • Release of DeFecT-FF, an online nanoHUB tool defect-ff with the following workflow: input bulk structure + list of defect candidates $\rightarrow$ generation of defect structures with symmetry breaking $\rightarrow$ MLFF optimization across five charge states $\rightarrow$ selection of the lowest-energy configurations $\rightarrow$ final HSE06+SOC (spin–orbit coupling) calculation.
2 Description of the DFT Datasets
The entire Cd/Zn–Te/Se defect chemical space is extraordinarily large, as summarized in the Supporting Information (SI), Table SI. Treating every symmetry-inequivalent site in a $3\times 3\times 3$ (216-atom) cubic zincblende supercell results in thousands of possible defect configurations per composition, driven by the combinatorial explosion of native vacancies, self-interstitials, anti-site substitutions, eight types of extrinsic defects (dopants—Cu, As, P, N, Sb, Bi; impurities—Cl, O), and all unordered defect complexes, as illustrated in Figure S1. Even when restricted to only neutral charge states, the number of required DFT calculations grows to nearly $\approx$ 0.9 million for the entire chemical space. Thankfully, the defect structures collected from published literature and our prior work Mannodi-ZB, Mannodi_Toriyama_2020, Mannodi-Kanakkithodi2022-ck, Rahman_Gollapalli already represent a physically meaningful and chemically rich subset of this much larger defect chemical space. These literature-curated PBE-optimized structures span multiple compositions, charge states, and defect chemistries, providing a robust foundation for training the initial GNN models. We then employed an active-learning workflow to selectively launch new PBE calculations in regions of the chemical space where the ML model exhibited high uncertainty or sparse representation. Details of this dataset-building strategy are provided in the SI.
A carefully selected subset of PBE-optimized structures was further used to perform hybrid HSE06 calculations after adjusting the defect supercell lattice parameters to values from HSE06 optimization, before volume-fixed relaxation, as summarized in the Table SII. The resulting HSE06 dataset captures more than half of the structural and chemical diversity present in the full PBE dataset, and the statistical distribution of this HSE subset is illustrated in Figure 1. Importantly, all HSE calculations preserved every ionic-relaxation snapshot—not only the final minimum-energy configuration—yielding thousands of intermediate structures with their corresponding energies, forces, and stresses. This provides a significantly enriched dataset that captures full relaxation pathways rather than ground-state endpoints alone. For a complete statistical overview of the PBE and HSE datasets across all charge states, readers are referred to Table SIII, Table SIV and Table SV. Additional dataset visualization is provided in Figure S2, Figure S3, and Figure S4. In summary, our data generation pipeline proceeds as follows: literature data collection $\rightarrow$ initial PBE GNN model training $\rightarrow$ active-learning–driven expansion of the PBE dataset $\rightarrow$ HSE06 refinement of a subset of the PBE dataset.
Figure  2: (a–c) Parity plots comparing crystal formation energies from DFT and MLFF predictions for three representative charge states: (a) q = +1, (b) q = 0 (neutral), and (c) q = –1. The MLFF accurately reproduces the DFT energies with small errors, as indicated by the RMSE values shown in each panel. (d–f) Defect formation energies computed using MLFF predictions for a subset of the defect configurations shown in panels (a–c), compared against values from full DFT. The MLFF defect energies were obtained by adding DFT reference energies and applying charge corrections to the MLFF-predicted total energies.
Figure  3: Workflow for accelerated defect predictions using the DeFecT-FF framework. An initial defect structure (example: Asi in a mixed “2Te–2Se” local environment) is constructed and passed through the ShakeNBreak snb_1 symmetry-breaking procedure to generate a diverse set of competing defect geometries. These distorted configurations are rapidly relaxed using the rigorously optimized machine-learned force field to identify the lowest-energy structure prior to high-fidelity DFT calculations. The optimized geometry is then used to perform static HSE+SOC calculation, yielding accurate defect formation energy diagrams.
3 MLFF models trained at hybrid functional accuracy
We initially trained models on the PBE dataset using the ALIGNN framework ALIGNN and then employed an active learning strategy to launch new DFT calculations by targeting regions of the chemical space with largest prediction uncertainty. We then refined a representative subset of the PBE-optimized structures using the HSE06 functional and used these data to train an M3GNet-based m3gnet MLFF, using DFT-derived configurations, energies, forces, and stresses. Readers are referred to the SI for additional details on the ALIGNN training procedure and the active learning workflow used in this work. These discussions are supplemented by a series of figures in the SI (Figures S5–S13), which collectively illustrate the active learning pipeline, model transferability, ALIGNN-based structural optimization, comparisons with other MLFF models, and MLFF models on PBE dataset.
HSE06 calculations were performed using $\Gamma$-point only Mosquera2, with a reduced plane-wave energy cutoff of 400 eV. The convergence thresholds for geometry optimization were set to $10^{-6}$ eV for energy and 0.01 eV Å-1 for forces. Figure 1 shows the statistics of the compiled HSE06 dataset in terms of the number of bulk Cd/Zn-Te/Se/S composition structures, and different types of defects. The HSE06 dataset represents 53.4% of the GGA dataset for bulk ($q=0$) structures, and 63.8%, 71.6%, 79.5%, 72.4%, and 65.8% for defect structures in charge states $q=+2$, $q=+1$, $q=0$, q=1q=-1, and q=2q=-2, respectively. Despite the smaller size, it remains well representative of the defect types and structural diversity of the entire chemical space. Violin plots showing the spread of the crystal formation energy (CFE, a per-atom energy difference between the crystal and its constituent atoms) values in the HSE06 dataset are presented in Figure S4.
Before training the MLFF models, we first evaluated the performance of state-of-the-art pretrained force-field frameworks, namely MACE mace, CHGNet chgnet and M3GNet m3gnet, by applying them directly to our PBE dataset. Although these models have demonstrated strong predictive capabilities on their training domains, they generalized poorly to the chemically diverse Cd/Zn–Te/Se/S systems investigated in this work as summarized in Table SVI. Root mean square error (RMSE) in CFE prediction ranged from 60 to 100 meV/atom, which it will turn out are much larger than errors from fine-tuned models. These shortcomings indicate that the pretrained models lack sufficient exposure to the chalcogenide defect chemical space considered here. Consequently, this motivated the development of a dataset-specific M3GNet-based MLFF trained on DFT-derived configurations, energies, forces, and stresses, enabling the level of accuracy required for reliable modeling of defect thermodynamics and structural relaxations.
Parity plots for M3GNet-MLFF models trained on the HSE06 dataset are pictured in Figure 2(a–c), respectively for charge states $q=+1$, $q=0$, and q=1q=-1. Models for the $q=+2$ and q=2q=-2 charge states are presented in Figure S15. Each parity plot compares the HSE06-computed CFE with MLFF-predicted values across different categories: bulk (pristine supercells without defects), and defects (bulk supercells containing a single defect or defect complex). Despite the reduced dataset size compared to the GGA dataset, the HSE-MLFF models achieve very good accuracy with low RMSE values across different structure types. The $q=0$ test set prediction RMSE ranges from 4.8 meV/atom for bulk structures to 7.8 meV/atom for defect structures. These errors are similar for $q=+1$ and q=1q=-1 defect structures and remain below 12 meV/atom for all cases, which is quite reasonable given the range of CFE values in the dataset. We also simulated a limited number of CdTe dislocation core structures Fatih-CdTe-GB, Jinglong and CdTe/ZnTe interface configurations with selected defects and included them in the training dataset. The model performance for these structures is shown in Figure S14 and Figure S15. These results are not discussed in detail in the main text because the number of data points corresponding to dislocation cores and interfaces is relatively small.
Figure S16 shows the MLFF predictions for neutral defect structures separately for single defects and defect complexes, revealing low RMSE values of 8.14 meV/atom and 9.23 meV/atom respectively. This suggests that the MLFF effectively captures both localized and collective defect relaxations, even for configurations with multiple defects. Even though MLFF prediction of crystal formation energy is highly accurate for all bulk and defect structures, a true evaluation of the prediction for defect configurations involves comparing defect formation energy (and defect level) predictions from MLFF and HSE06. We used the MLFF models to optimize a selected set of defects structures. For each defect, the MLFF was used to compute the total energy entering the standard defect formation energy expression:
$\Delta E_{f}(D^{q})=E_{tot}(D^{q})-E_{tot}(bulk)+\sum_{i}n_{i}\mu_{i}+q\bigl(E_{F}+E_{VBM}\bigr)+E_{corr}$
Here, $E_{tot}(D^{q})$ and $E_{tot}(bulk)$ are the total energies of the defect and pristine supercells respectively, $n_{i}$ and $\mu_{i}$ denote the stoichiometric changes and elemental chemical potentials, $q$ is the charge state, $E_{VBM}$ is the valence band maximum energy, $E_{F}$ is the Fermi level through the band gap, and $E_{corr}$ is the correction energy Freysoldt2014-en which accounts for spurious electrostatic interactions arising from the periodic repetition of charged defects and the compensating background charge in finite supercells. In this work, $E_{corr}$ is evaluated using the Freysoldt  Freysoldt2014-en charge correction scheme, which separates long-range Coulomb interactions from short-range defect-induced potentials and aligns the electrostatic potential between defect and bulk calculations Freysoldt2014-en. Additional methodological details and convergence tests for the charge correction are provided in the SI. Importantly, the reference energy $\mu_{i}$, $E_{VBM}$ and $E_{corr}$ were evaluated from DFT and added directly to the MLFF-derived bulk and defect energies.
Figure 2(d–f) present parity plots for defect formation energy corresponding to $q=+1$ (at EF=0), $q=0$, and q=1q=-1 (at EF=0). Across the entire validation set, the RMSE in defect formation energies obtained using MLFF-optimized geometries remains below $0.20$ eV, demonstrating that the MLFFs are sufficiently accurate in capturing both structural and energetic trends for charged and neutral defects. To further assess the influence of charge corrections on MLFF-derived defect energetics, we compared three approaches: (i) using MLFF-predicted defect formation energies without any charge correction; (ii) using MLFF defect formation energies corrected using a simple average offset of $0.20$ eV for $q=+2$, $0.10$ eV for $q=+1$, $0.10$ eV for q=1q=-1, and $0.20$ eV for $q=+2$ defects; and (iii) adding known charge correction energies from DFT to MLFF-predicted defect formation energies (Figure 2(d–f)).
The comparison reveals that while uncorrected MLFF prediction shows errors close to 0.3 eV for all charge states, applying the average offset brings this error down closer to 0.2 eV which is similar to the error from adding known correction values, as listed in Table SVII. Parity plots comparing MLFF defect formation energy with DFT values for all three approaches are shown in Figure S17. Figure S18 compares defect charge transition levels predicted by the MLFF with DFT reference values. Applying an average charge correction value leads to close agreement between MLFF- and DFT-predicted defect transition levels, with RMSE values of 0.25 eV, 0.23 eV, 0.22 eV, and 0.27 eV for the $\epsilon(+2/+1)$, $\epsilon(+1/0)$, $\epsilon(0/-1)$, and $\epsilon(-1/-2)$ transitions, respectively, whereas the DFT-based charge correction values are 0.22 eV, 0.21 eV, 0.20 eV, and 0.24 eV.
To assess the reliability of our MLFF models relative to full DFT, we randomly selected 100 representative bulk and defect configurations and relaxed each structure using both methods. The DFT- and MLFF-optimized geometries were then compared using SOAP descriptors dscribe_1, dscribe_2, which provide a rotationally invariant fingerprint of the atomic environments. These high-dimensional descriptors were projected into a two-dimensional PCA PCA (principle components analysis) space, allowing direct visualization of structural similarity. In Figure S19, each DFT structure (blue) is paired with its corresponding MLFF structure (orange), with a connecting line indicating the degree of agreement. The consistently short line segments demonstrate that the MLFF reproduces DFT relaxation behavior with high accuracy. Some representative examples of MLFF-based defect structure optimizations are shown in Figure S11(d–f) and Figure S14(d–f) of the SI.
Figure  4: Benchmarking DeFecT-$F$ for a selected defect complex in CdSe0.12Te0.88. (a) Visualization of the AsSe + ClSe complex in the CdSe0.12Te0.88 alloy supercell. (b) Total energy relaxation profiles for different charge states, comparing the converged DFT energies with the DeFecT-FF–relaxed energies. (c) Defect formation energies under Cd-rich conditions for charge states +2+2 to 2-2, showing close agreement between DFT and DeFecT-FF predictions. (d) Computational cost, measured in core-hours, highlighting the significant reduction in wall-time achieved when using DeFecT-FF instead of full DFT relaxations.
Figure  5: Defect formation energy diagrams for AsX, ClX, and AsX+ClX defects in (a) CdTe and (b) CdSe0.25Te0.75, under Cd-rich conditions; X = Te or Se. (c) Defect charge transition levels for AsX, ClX, and AsX+ClX, computed for different CdSexTe1-x compositions ($x=0.0$, 0.06, 0.12, 0.25). Blue lines indicate the AsX (0/–1) acceptor level, red lines show the ClX (+1/0) donor level, and purple lines show the AsX+ClX (0/–1) level. For each compound, the VBM is placed at EF = 0 eV and the CBM is placed at the value of the computed band gap. All results are from HSE06+SOC calculations performed after DeFecT-$F$ optimization.
4 New Predictions with the MLFF Models: Case Studies of Important Defects
The MLFF models can now be used to optimize any given defect configuration in different charge states with near–hybrid-functional accuracy. Following MLFF optimization, final HSE06+SOC single-point calculations must be performed to obtain reliable defect formation energy diagrams. Figure 3 illustrates the overall workflow of the DeFecT-FF framework defect-ff in determining the lowest energy symmetry-broken defect configuration followed by a high-fidelity understanding of the defect thermodynamics. Figure S20 shows the workflow of the DeFecT-FF defect-ff web tool we created on the nanoHUB platform to enable efficient creation of defect structures, MLFF optimization, and visualization of defect formation energy diagrams. In the next subsections, we present a few case studies demonstrating the application of this workflow to determine the relative stability and charge transition levels of important defects in selected Cd/Zn-Te/Se/S compositions which were not entirely part of the MLFF training dataset.
4.1 As+Cl Defect Complex in CdSe0.12Te0.88
As is a commonly used p-type dopant in Se-alloyed CdTe and Cl is a common impurity arising from CdCl2 treatment, which makes it important to investigate defect complexes of As and Cl in representative CdSeTe compositions. We applied the DeFecT-FF framework to simulate the AsSe + ClSe substitutional defect complex in the compound CdSe0.12Te0.88, across five possible charge states. An example configuration is illustrated in Figure 4(a); Se alloying starting from a 216-atom CdTe cubic supercell is first simulated using the special quasirandom structures (SQS) approach SQS, following which As and Cl substitution is incorporated in multiple possible locations to eventually yield the preferred combination. For each charge state between $q=+2$ and q=2q=-2, ten symmetry-broken initial structures were generated using the ShakeNBreak protocol snb_1, enabling exploration of a diverse set of competing local geometries. Each of these structures was relaxed using DeFecT-FF to identify the lowest-energy configuration prior to high-fidelity DFT refinement.
The total energy relaxation profiles for different charge states are shown in Figure 4(b). The DeFecT-FF structural optimizations converge smoothly and yield geometries that lie very close to those produced by DFT calculations. After DeFecT-FF relaxation, a single-shot HSE06+SOC calculation is performed on the predicted lowest-energy geometry for each charge state to accurately compute defect formation energies. The resulting defect formation energies (at EF=0) under Cd-rich conditions from DFT and DeFecT-FF are compared in Figure 4(c). Across all charge states, the agreement is excellent, with typical deviations well below 0.1–0.2 eV.
Thus, the DeFecT-FF geometries provide a sufficiently accurate structural foundation for defect thermodynamics for complexes in an alloyed CdSeTe composition. The computational savings are substantial: a single HSE06 relaxation of a charged defect in a $\times3$$\times$3 supercell requires approximately 512 cores multiplied by 8 to 9 hours per configuration, corresponding to a total of nearly 4096 core-hours. In contrast, the DeFecT-FF relaxations require only about $(2/60)\times 16$ core-hours per configuration, which is approximately 0.5 core-hours. The speedup therefore exceeds four orders of magnitude, as presented in Figure 4(d). We expect these trends to hold for all types of defect complexes in alloyed $\times3$$\times$3 supercells and there is confidence in DeFecT-FF reaching close agreement with hybrid DFT at a fraction of the cost.
4.2 As and Cl Defects Across CdSexTe1-x Compositions
Next, we simulated multiple substitutional defects of As and Cl (including complexes) in $3\times 3\times 3$ supercells of a series of CdSexTe1-x compositions (x=0, 0.06, 0.12, 0.25x=0,\,0.06,\,0.12,\,0.25). Using the Doped doped package, we introduced defects AsTe, AsSe, ClTe, ClSe and the AsX+ClX double defect complexes (where $X$ denotes the preferred anion site, Te or Se). Symmetry‐breaking operations were then applied via the ShakeNBreak protocol, enabling the sampling of a diverse set of competing configurations. Hundreds of structures for these substitutional defects across the CdSexTe1-x compounds were relaxed with the DeFecT-FF models for different charge states until the maximum force fell below $10^{-5}$ eV/Å. Finally, single‐shot HSE06+SOC calculations were performed to obtain accurate defect formation energy.
The band gaps computed using HSE06+SOC (with a modified mixing parameter of $\alpha$=0.31) for CdTe, CdSe0.06Te0.94, CdSe0.12Te0.88, and CdSe0.25Te0.75 are respectively 1.5 eV, 1.41 eV, 138 eV, and 1.30 eV; these values are used to place the EF bounds for the defect formation energy diagrams. The VBM for each composition was obtained from the bulk calculation at the HSE+SOC level using a $2\times 2\times 2$ $k$-mesh for the corresponding 216-atom $3\times 3\times 3$ supercell. The charge-dependent defect formation energies additionally yield the charge transition levels as described below:
$\varepsilon(q/q^{\prime})=\frac{\Delta E_{f}(D^{q};E_{F}=0)-\Delta E_{f}(D^{q^{\prime}};E_{F}=0)}{q^{\prime}-q}$
This transition level marks the EF position at which charge states $q$ and $q^{\prime}$ are in equilibrium. Figure 5 presents the defect formation energy diagrams and relevant transition levels for AsX, ClX, and AsX+ClX defects across the CdSexTe1-x series, with EVBM set to 0 eV; X represents either Te or Se. Incorporation of Se is observed to deepen the AsX 0/-1 acceptor level despite the band gap going down from CdTe to CdSe0.25Te0.75, in agreement with recent experimental studies Nardone. The ClX +1/0 donor level remains deep in the band gap in all cases, around 1 eV from the VBM, while the AsX+ClX defect complex, interestingly, creates a 0/-1 acceptor level closer to the conduction band edge which becomes shallower with more Se content due to the lowering of the CBM. The defect energy diagrams in Figure 5(a-b) show the prevalence of the neutral state for the defect complex in the band gap, while AsX and ClX respectively create low energy acceptor and donor defects which pin the equilibrium EF (obtained by applying charge-neutrality conditions) around the middle of the band gap.
4.3 Native Defects and Nitrogen Impurities in ZnTe
Motivated by experimental evidence from X-ray photoelectron spectroscopy (XPS) f__a__stevie_72130f8a, janet_mahoney_e1a9d305, artur_born_2c82645b, gustavo_lanza_7007a95e, hui_chen_dca65eb5, haipeng_xie_92ab6279 indicating N incorporation in ZnTe ji_ha_lee_5db163c4, lijuan_zhao_6c062f36, yingrui_li_bb64582f, titao_li_7535f555, daniele_dragoni_41c08f2d, eduardo_men_ndez_proupin_dae1679b, ethan_berger_ad4781d2, we employed the DeFecT-FF workflow to systematically investigate both native point defects and N-related defects in ZnTe. A $3\times 3\times 3$ ZnTe supercell (cubic zinblende phase) was first fully relaxed using the HSE06 functional prior to defect introduction; its band gap was computed to be 2.2 eV from HSE06+SOC. The defect set included vacancies, interstitials, and antisite defects (VZn, VTe, Zni, Tei, ZnTe, TeZn), as well as the following N defects: Ni, NTe, Ni+Ni, and NTe+Ni. To ensure thorough exploration of the potential energy landscape, we applied ShakeNBreak snb_1, snb_2 to induce perturbations, enabling the sampling of a diverse set of competing configurations. Among the hundreds of N-related configurations evaluated, the Ni+Ni defect complex emerged as the most energetically favorable, a finding further validated through additional HSE06+SOC calculations. Figure 6(a) illustrates the DeFecT-FF structural optimization and energy convergence for Ni+Ni in ZnTe, and Figure 6(b) presents the HSE06+SOC computed defect formation energy diagram for different N-related defects in ZnTe.
Figure  6: (a) Energy as a function of optimization steps during DeFecT-$F$ relaxation of a ZnTe supercell with the double N interstitial (Ni+Ni) defect complex. The inset shows the relaxed configuration with N atoms in red. (b) Defect formation energies in ZnTe under Te-rich conditions computed using HSE06+SOC on top of the HSE-MLFF optimized configurations.
Figure  7: (a) Total energy as a function of simulation time for a microcanonical (NVE) molecular dynamics trajectory of the AsTe defect in CdTe, demonstrating excellent energy conservation over a 100 ps run. The inset shows the atomic structure before (initial) and after (final) MD simulation, with the defect site highlighted. (b) Normalized vibrational density of states (VDOS) obtained from the velocity autocorrelation function of the same trajectory, revealing dominant low-frequency vibrational modes below 4 THz that are characteristic of defect-localized and heavy-atom vibrations.
4.4 Simulating Defects at Finite Temperature
Defect formation energies in semiconductors are typically evaluated at $T=0$ K using DFT, neglecting vibrational entropy effects that can become important under realistic growth and operating conditions. While a full finite-temperature free-energy treatment is beyond the scope of this work, we demonstrate that DeFecT-FF enables finite-temperature molecular dynamics (MD) MD_1, MD_2 simulations that provide a physically grounded pathway toward incorporating such effects in future studies. As a representative example, we consider the AsTe defect in CdTe. The defect structure was first relaxed using DeFecT-FF, followed by finite-temperature microcanonical (NVE) ensemble MD simulations for the neutral charge state ($q=0$) at $T=300$ K using a 1 fs time step and a total simulation length of 100,000 steps MD_3, MD_4, MD_5. Figure 7(a) shows the energy as a function of simulation time, demonstrating excellent energy conservation over the 100 ps trajectory and confirming the numerical stability of DeFecT-FF.
Beyond validating stability, finite-temperature MD offers key physical advantages over static relaxations by enabling symmetry-breaking driven by thermal fluctuations and realistic atomic motion. This allows the system to explore the local configurational space and access dynamically stabilized or lower energy defect structures that may not be achieved at 0 K. Such MD-relaxed structures provide physically meaningful inputs for defect formation energy calculations and establish a foundation for future investigations of vibrational entropy and finite-temperature defect thermodynamics using the same framework.
During the MD simulations, atomic velocities $v_{i}(t)$ were recorded at each time step, where $i$ labels atoms in the supercell and $N$ is the total number of atoms. From these trajectories, we computed the velocity autocorrelation function as follows Rahman_Sun:
$C_{vv}(t)=\frac{1}{3N}\sum_{i=1}^{N}\langle v_{i}(0)\cdot v_{i}(t)\rangle,$ (1)
and obtained the vibrational density of states (VDOS), $g(\omega)$, using Fourier transformation:
$g(\omega)=\int_{0}^{\infty}C_{vv}(t)\,e^{-i\omega t}\,dt,$ (2)
Here, $\omega=2\pi f$ is the vibrational angular frequency. Figure 7(b) shows the normalized VDOS derived from the MD trajectory. The dominant low-frequency modes below approximately 4 THz are characteristic of defect-localized vibrations and reflect vibrational softening induced by the AsTe defect and the presence of heavy atoms in CdTe. The VDOS obtained from finite-temperature MD naturally incorporates anharmonic effects and provides direct access to vibrational entropy contributions Freysoldt2014-en which will be addressed in follow-up contributions from our group.
5 Conclusions
CdSexTe1-x solar cells are fundamentally constrained by defect physics: deep-level nonradiative centers from native defects and impurities limit open-circuit voltage, dopants such as Cu and As often lead to unhelpful complexes, and extended defects at interfaces and grain boundaries act as sinks for charge and sites for defect clustering. While hybrid-functional DFT remains the gold standard for resolving these mechanisms, its cost prevents exhaustive exploration across alloy compositions, charge states, and structural motifs. To overcome these barriers, we introduced the DeFecT-FF framework in this work, a crystal graph-based active learning-driven MLFF model trained on data from both semi-local GGA and hybrid HSE06 calculations for thousands of charged and neutral structures spanning the Cd/Zn–S/Se/Te chemical space, with a wide variety of native and extrinsic defects and defect complexes considered. DeFecT-FF predicts energies and forces across charge states, enabling rapid geometry optimization and defect formation energy evaluation. We demonstrated the utility of these models by identifying low energy configurations of device-relevant defects and performing HSE06+SOC calculations to understand their energetics and defect levels.
In practice, the DeFecT-FF framework reduces single defect optimization time from at least $\sim$8–9 h (HSE06) to $\sim$1–2 min while retaining near-DFT accuracy, transforming comprehensive, composition- and charge-resolved defect surveys from intractable to routine. We have deployed this framework as part of a Jupyter notebook-based nanoHUB tool which will allow users to upload CIF files of Cd/Zn-Te/Se/S structures, auto-generate relevant defects or complexes, and compute their defect formation energies as functions of Fermi level and chemical potentials conditions, bypassing expensive first principles workflows. Together, these advances provide a scalable, charge-aware pathway to map defect landscapes in chemistries relevant to CdSeTe solar cell devices and beyond, accelerating the dopant/process optimization and ultimately closing the voltage deficit in this important thin-film photovoltaic platform.
Conflicts of Interest
There are no conflicts to declare.
Data Availability
All raw and processed data supporting this work, namely atomic structures, total energies, forces, and stresses, and the trained MLFF models used in this study, are publicly accessible via a nanoHUB web application at https://nanohub.org/tools/cadetff. This tool enables community reuse and verification through a point-and-click interface. Within the same tool, users can:
  • Provide inputs: Upload a crystallographic file (e.g., CIF or POSCAR), select native or extrinsic defect types (vacancy, interstitial, substitutional, or complexes).
  • Run computations: Perform MLFF-based geometry relaxation in minutes and obtain optimized energies to identify lowest energy configurations.
  • Retrieve outputs: Download relaxed structures, per-defect energies, and tabulated CSV summaries, along with run logs for reproducibility.
  • Intended scope: Rapid screening and hypothesis testing; device-critical cases should be validated with targeted HSE06(+SOC) calculations.
Funding Statement
This material is based upon work supported by the U.S. Department of Energy’s Office of Energy Efficiency and Renewable Energy (EERE) under the Solar Energy Technology Office (SETO) Award Number DE-0009332. Funding for this work was also provided by the Alliance for Sustainable Energy, LLC, Managing and Operating Contractor for the National Renewable Energy Laboratory for the U.S. DOE, and was supported in part by EERE under SETO Award Number 37989. A.M.K. additionally acknowledges support from Argonne National Laboratory under sub-contracts 21090590 and 22057223, from DOE EERE. This research used resources from the the Center for Nanoscale Materials (CNM) at Argonne National Laboratory. Work performed at the CNM, a U.S. Department of Energy Office of Science User Facility, was supported by the U.S. DOE, Office of Basic Energy Sciences, under Contract No. DE-AC02-06CH11357. This work also utilized the Anvil cluster at Purdue through allocation MAT230030 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
Acknowledgments
The authors would like to acknowledge discussions with Dr. Mariana Bertoni at Arizona State University, Dr. Yanfa Yan at University of Toledo, and researchers at the National Renewable Energy Laboratory. We also acknowledge the Rosen Center for Advanced Computing (RCAC) clusters at Purdue University for further computational support.
Author Contributions Statement
A.M.-K. conceived and planned the research project and procured research funding. DFT computations and MLFF training tasks were performed by M.H.R. and M.B. M.H.R. took the lead on writing; M.B. and A.M.-K. contributed in editing and shaping the manuscript.
Supplemental material to "DeFecT‑FF: Accelerated Modeling of Defects in Cd/Zn-Te/Se/S Compounds Combining High-Throughput DFT and Machine Learning Force Fields"
Md Habibur Rahman,1) Maitreyo Biswas,1) and Arun Mannodi-Kanakkithodi 1, a)
1School of Materials Engineering, Purdue University, West Lafayette, Indiana 47907, USA
footnotetext: aamannodi@purdue.edu
Table SI: Enumeration of defects across binary and alloyed Cd/Zn–Te/Se compositions. Binary compounds (CdTe, CdSe, ZnTe) contain two atomic species and therefore host 6 native defect types. Alloyed compounds contain three distinct ionic species, yielding 12 native defect types. Extrinsic defect species (Cu; As, P, N, Sb, Bi; Cl, O) contribute three defect types each (interstitial, cation substitution, anion substitution). For mixed compositions, we enforce a minimum of 10 non-equivalent positions per defect type, even through thousands of distinct symmetry-inequivalent positions exist (e.g., Asi in Cd–Se–Te). Each defect configuration is additionally perturbed using ShakeNBreak snb_2 to generate at least 15 symmetry-broken initial structures.
Compound Native Extrinsic Native + Extrinsic Double complexes Total unique defects Non-eq sites ($\times$10) Perturbed structures ($\times$15)
CdTe 6 24 30 435 465 - 6,975
CdSe 6 24 30 435 465 - 6,975
ZnTe 6 24 30 435 465 - 6,975
CdSe0.25Te0.75 12 32 44 946 990 9,900 148,500
CdSe0.50Te0.50 12 32 44 946 990 9,900 148,500
CdSe0.75Te0.25 12 32 44 946 990 9,900 148,500
Cd0.25Zn0.75Te 12 32 44 946 990 9,900 148,500
Cd0.50Zn0.50Te 12 32 44 946 990 9,900 148,500
Cd0.75Zn0.25Te 12 32 44 946 990 9,900 148,500
Composition PBE (Å) HSE06 (Å)
CdSexTe1-x
CdTe ($x=0.00$) 19.88 19.72
CdSe0.25Te0.75 19.57 19.34
CdSe0.50Te0.50 19.25 19.02
CdSe0.75Te0.25 18.94 18.71
CdSe ($x=1.00$) 18.63 18.45
CdxZn1-xTe
CdTe ($x=1.00$) 19.88 19.72
Cd0.75Zn0.25Te 19.56 19.35
Cd0.50Zn0.50Te 19.21 19.01
Cd0.25Zn0.75Te 18.85 18.72
ZnTe ($x=0.00$) 18.52 18.42
Table SII: Optimized supercell lattice parameters (in Å) for CdSexTe1-x and CdxZn1-xTe compounds computed from PBE and HSE06.
Figure  S1: Schematic illustration of the combinatorial expansion of the full Cd/Zn–Te/Se defect chemical space across binary and ternary compositions. Each ternary alloy (Cd–Zn–Te and Cd–Se–Te) hosts 44 unique single-defect configurations arising from vacancies, antisites, interstitials, and eight extrinsic defects (N, P, As, Sb, Bi, Cu, O, Cl). These yield $^{44}C_{2}=946$ unique defect complexes. Combining all single defects and complexes gives 990 distinct defect types per composition, which are further expanded by sampling at least 10 non-equivalent lattice positions per defect and generating 15 ShakeNBreak snb_2 symmetry-broken perturbations per configuration, resulting in approximately $\sim 1.495\times 10^{5}$ total structures per ternary alloy. For the binary compound CdTe (and similarly CdSe and ZnTe), 30 unique single defects generate $^{30}C_{2}=435$ complexes, yielding unique 465 defects, each further expanded through 15 perturbed initial structures to give 6,975 configurations.
Detailed Description of the DFT Datasets
The full DFT dataset used in this work includes two level of theories: a broad GGA–PBE dataset containing tens of thousands of bulk and defect configurations, and subset of PBE data refinement using the HSE06 hybrid functional. The majority of the PBE dataset was collected from previously published works from our group covering defects in a wide range of II–VI semiconductors Mannodi-ZB, Mannodi_Toriyama_2020, Mannodi-Kanakkithodi2022-ck, Rahman_Gollapalli. These structures were combined with a significant number of new PBE calculations, particularly for alloy compositions and defect configurations that were not part of prior studies. Our chemical space spans binary Cd/Zn–Te/Se/S compounds, selected ternaries, and several quaternary alloys. Table SIII lists the full set of bulk compounds included in this study. Each composition is simulated using both $2\times 2\times 2$ (64-atom) zincblende supercells generated using the special quasirandom structure (SQS) approach SQS. For each composition, the mixing fraction $x$ is systematically varied in increments of $0.125$ (i.e., $n/8$ for $n\in\{0,\dots,8\}$), yielding a total of 81 unique compositions Mannodi-ZB.
Table SIII: Summary of bulk compounds included in the PBE/HSE dataset. The full set spans binary, ternary, and quaternary compositions in the Cd/Zn–Te/Se/S chemical space.
Category Compounds
Binary CdTe, CdSe, CdS, ZnTe, ZnSe, ZnS
Ternary alloys CdSexTe1-x, CdSxSe1-x, ZnSexTe1-x, ZnSxSe1-x
Quaternary alloys CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x
Across all compositions and supercell sizes, the PBE dataset contains more than 10,000 bulk structures, including not only the fully relaxed configurations but also all intermediate snapshots along the geometry-optimization trajectories. Each structure is labeled by its crystal formation energy (CFE), defined for a general composition CdaSebTec as:
$CFE=\frac{E(\text{Cd}_{a}\text{Se}_{b}\text{Te}_{c})-aE(\text{Cd})-bE(\text{Se})-cE(\text{Te})}{N_{\text{atoms}}}.$ (3)
Here, $E(\text{Cd${}_{a}$Se${}_{b}$Te${}_{c}$})$ is the total DFT energy of the supercell containing a atoms of Cd, b atoms of Se, and c atoms of Te that is necessary to simulate the CdSexTe1-x composition. $E(\text{Cd})$, $E(\text{Se})$, and $E(\text{Te})$ are respectively the per-atom energies of Cd, Se, and Te in their known elemental standard states, and $N_{\text{atoms}}$ = a+b+c is the total number of atoms in the supercell.
The defect dataset contains native vacancies, interstitials, antisites, extrinsic interstitials and substitutional dopants (Cu, As, P, N, Sb, Cl, O), and selected complexes across multiple compositions. All defect calculations were performed in either $2\times 2\times 2$ or $3\times 3\times 3$ supercells, and multiple charge states were considered. Additionally, we incorporated CdTe dislocation–core structural models derived from prior STEM-based work Fatih-CdTe-GB. These structures correspond to a 4.8 low-angle $\parallel(110)$ tilt boundary, in which the extended defect is represented by a periodic array of Lomer-type dislocation cores. Two experimentally observed core reconstructions were simulated: a Type-I core, where the absence of a central Cd column produces a Te-terminated dislocation core with broken Te–Cd bonds that generate deep mid-gap states; and a Type-II core, where a restored central Cd column leads to a distinct mixed Cd–Te coordination environment and an electronic state. Together, these two structures capture the essential bonding rearrangements, dangling-bond motifs, and charge perturbations intrinsic to realistic CdTe dislocation cores identified through atomic-resolution STEM imaging Fatih-CdTe-GB. We also constructed a CdTe–ZnTe heterointerface by merging fully relaxed $\times3$$\times$3 supercells of CdTe and ZnTe. Small in-plane strains were applied to lattice-match the two materials, followed by volume-conserving structural relaxation to obtain a physically consistent interface. For all new defect calculations in the $\times3$$\times$3 binary, alloyed, and interface the Doped package doped was used to generate symmetry-broken initial configurations via bond distortions and atomic rattling. In the unified dataset, the label bulk denotes pristine supercells; defect refers to supercells containing vacancies, interstitials, antisites, dopants, or multi-defect complexes; interface corresponds to CdTe–ZnTe heterostructures containing defects at or near the interfacial region; and dislocation_core designates structures based on the Type-I and Type-II STEM-derived dislocation-core models, with defects introduced within their reconstructed core environments. The complete PBE dataset distributions used for ALIGNN training are reported in Figure S2, Figure S3, and Table SIV.
A curated subset of bulk and defect structures from the PBE dataset was re-optimized using the HSE06 hybrid functional with a mixing parameter of $\alpha=0.25$. Due to high computational costs, HSE06 calculations were performed only for representative configurations spanning different compositions, and chemical environments. For defect calculations, lattice parameters were updated to the HSE06-optimized bulk volume before relaxation. All hybrid-functional calculations used $\Gamma$-point sampling and reduced plane-wave cutoffs appropriate for large supercells. The distributions of the complete HSE dataset are presented in Figure S4 and Table SV. This dataset forms the basis for the MLFF ALIGNN models used in this work.
Dataset Supercell Size Data Points Bulk dataset from Cd/Zn-Te/Se/S binary, ternary and quaternary alloys $\times2$$\times$2 10080 Bulk dataset from CdSexTe1-x alloys $\times3$$\times$3 26 Defect dataset from 6 Cd/Zn-Te/Se/S binary compounds $\times2$$\times$2 7302 (q=+2), 6201 (q=+1), 8203 (q=0), 6361 (q=-1), 7689 (q=-2) Defect dataset from CdSexTe1-x and CdxZn1-xTe alloys $\times3$$\times$3 594 (q=+2), 533 (q=+1), 643 (q=0), 514 (q=-1), 560 (q=-2) Defect dataset from CdTe/ZnTe interface $\times3$$\times$6 375 (q=+2), 380 (q=+1), 401 (q=0), 381 (q=-1), 330 (q=-2) Defect dataset from CdTe dislocation core 210 (q=+2), 223 (q=+1), 263 (q=0), 220 (q=-1), 231 (q=-2)
Table SIV: Number of data points (or structures) in the GGA-PBE dataset corresponding to different types of bulk or defect configurations, supercell sizes, and charge states.
Dataset Supercell Size Data Points Bulk dataset from Cd/Zn-Te/Se/S binary compounds and alloys $\times2$$\times$2 5400 Defect dataset from 6 binary compounds $\times2$$\times$2 4302 (q=+2), 4201 (q=+1), 6203 (q=0), 4361 (q=-1), 4689 (q=-2) Defect dataset from CdSexTe1-x and CdxZn1-xTe alloys $\times3$$\times$3 371 (q=+2), 333 (q=+1), 402 (q=0), 321 (q=-1), 350 (q=-2)
Table SV: Number of data points (or structures) in the HSE06 dataset corresponding to different types of bulk or defect configurations, supercell sizes, and charge states.
Figure  S2: Violin distributions of crystal formation energy per atom (meV) for charge states +2+2, +1+1, 0 (neutral), 1-1, and 2-2 in the GGA-PBE dataset.
Figure  S3: Statistics of the GGA-PBE dataset: (a) Number of bulk configurations corresponding to CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x, ZnSxSe1-x, and ZnSexTe1-x compositions. (b) Number of neutral defect configurations in CdS, CdSe, CdTe, ZnS, ZnSe, ZnTe, and different CdSexTe1-x compositions, with the inset showing the distribution of vacancy, substitutional, and interstitial defects across the dataset. (c) Bar chart showing the total number of bulk, defect, interface, and dislocation core (DC) configurations in the dataset. (d) Violin plots showing the distribution of crystal formation energy across all the defect, bulk, interface and DC structures. Inside each violin, a mini box plot shows the median (central line), quartile, and range (whiskers) (e, f, g) Distribution of defect configurations for different charge states (q = +2, +1, 0, -1, and -2) in bulk defect, interface, and DC configurations, respectively.
Figure  S4: Violin distributions of crystal formation energy per atom (meV) for charge states +2+2, +1+1, 0 (neutral), 1-1, and 2-2 in the HSE06 dataset.
Figure  S5: Parity plots for ALIGNN models trained on the GGA dataset using (a) only bulk structures, (b) only defect structures, with predictions distinguished in terms of type of defect, and (c) both bulk and defect structures. (d) Active learning (AL) workflow implemented in this work: standard deviation of ALIGNN-predicted formation energy of novel defects is used to determine acquisition functions and identify the next set of DFT simulations to run. The ALIGNN model is then retrained with the new data and predictions are made for the remaining set of unexplored defects. (e) An ALIGNN vs DFT parity plot showing standard deviation in test set prediction across 100 separate models; these error bars are used to calculate the acquisition functions in the AL workflow.
Figure  S6: Parity plots comparing ALIGNN predictions to DFT calculations under different training conditions: (a) ALIGNN (trained solely on bulk structures) predictions vs. DFT calculations for the defect dataset. (b) ALIGNN (trained exclusively on interstitial defect structures) predictions vs. DFT calculations for the vacancy dataset. (c) ALIGNN (trained only on $\times2$$\times$2 supercell bulk and defect structures) predictions vs. DFT calculations for defects in a $\times3$$\times$3 supercell.
Graph neural network models for direct prediction of crystal formation energy
ALIGNN was developed by Choudhary et al.ALIGNN and considers both two-body (bond lengths) and three-body interactions (bond angles). ALIGNN leverages both graph convolution layers and line graph convolution layers to capture short-range and long-range correlations in the crystal. For training the ALIGNN models, the learning rate was set to 0.001, an AdamW optimizer was used to update the weights and biases of the model, 4 graph convolution layers and 4 line graph layers were implemented, the cutoff radius was set to 6 Å with 12 nearest neighbors to create the crystal graph, and models were trained up to 90 epochs with a batch size of 8. We experimented with different training-validation-test splits of the dataset and found that the 60:20:20 ratio works the best.
ALIGNN models were trained to predict the CFE (PBE) from any given bulk or defect crystal structure, using only the neutral charge state structures at this stage. Parity plots capturing the performance of the optimized models are presented in Figure S5, in terms of ALIGNN-predicted CFE vs DFT-computed CFE for only the test set data points. Models pictured in Figure S5(a-c) are respectively trained only on bulk structures, only on defect structures, and on both bulk and defect structures; this distinction is made to understand how sensitive the models are to different types of configuration. As shown in Figure S5(a), the ALIGNN model for bulk structures alone shows a test prediction root mean squared error (RMSE) of 1.43 meV/atom, and this error remains practically unchanged for the combined model in Figure S5(c). For the defect-only ALIGNN model in Figure S5(b), test RMSE ranges from 3.09 meV/atom for interstitial defects to 4.87 meV/atom for substitutional defects to 8.36 meV/atom for vacancy defects. Each of these defect prediction errors comes down for the combined data model, proving the value of increasing the size and chemical and structural diversity of the training dataset.
The training dataset contains a larger number of interstitial defects, followed by substitutional and vacancy defects: this is mostly a consequence of there being a lot more options for intersitial and substitutional defects in terms of extrinsic species from across the periodic table, and also the longer time it takes for DFT optimization of these defect structures, leading to more intermediate geometries. This results in the comparatively higher RMSE for CFE prediction of vacancy defects (6.47 meV/atom) than substitutional (3.84 meV/atom) and interstitial (2.04 meV/atom) defects. The ALIGNN predictions for all types of structures are highly accurate with vanishingly small errors considering the total range of CFE values. GNN models for defect predictions reported in the recent literature Mosquera2, Witman_2023, Rahman_Gollapalli primarily focused on sampling defect configurations to train surrogate models, while we have adopted the approach of combining bulk and defect structures which enhances the overall generalizability and accuracy of the models.
Figure  S7: (a) Violin plot showing the mean crystal formation energy (CFE) predicted by the initial ALIGNN-1 models, averaged across 100 models. (b) Defects with maximum uncertainties identified through ALIGNN-1 models. (c) Comparison of ALIGNN-1 predictions vs DFT calculations for 200 selected defects. (d) Violin plot of mean CFE from ALIGNN-2 models. (e) Defects with maximum uncertainties identified through ALIGNN-2 models. (f) Comparison of ALIGNN 2 predictions vs DFT calculations for the 200 selected defects.
Although state-of-the-art GNNs are very robust and powerful for modeling complex relationships within atomic structures Choudhary_DeCost, ALIGNN, Choudhary, MEGNET, CGCNN, Cheng_Dong_2021, Lee_Asahi, Fung, their transferability beyond the trained dataset remains questionable. To evaluate this, we performed the following series of tests:
  1. 1.
    An ALIGNN model was trained purely on bulk structures and then used to predict the CFE of defect structures.
  2. 2.
    An ALIGNN model was trained only on interstitial defect structures and used to predict the CFE for vacancy and substitutional defects.
  3. 3.
    An ALIGNN model trained exclusively on defects in $\times2$$\times$2 supercell structures and then used to predict the CFE of $\times3$$\times$3 supercell defect structures.
ALIGNN shows poor transferability when trained only on bulk data and used to predict for defects; as shown in Figure S6(a), the prediction RMSEs range from $\sim$ 36 meV/atom for vacancies to >> 50 meV/atom for interstitial and substitutional defects. Since this model has not been exposed to specific configurations such as atomic relaxation around defect sites, it is unable to predict as accurately for defect structures as it does for bulk. Figure S6(a) also shows that ALIGNN uniformly under-predicts the CFE of all defect configurations, which could be attributed to the lower average CFE values in the bulk dataset compared to defects, as illustrated by the violin plot in Figure S3(e). The model trained only on interstitial defects does a reasonable job for vacancy defects, but the RMSE values are larger than from the models in Figure S5 and there are some very clear outliers, as pictured in Figure S6(b). Lastly, when the model is trained on only $\times2$$\times$2 supercell structures and used to predict for $\times3$$\times$3 supercells, the predictions show good accuracy but a slight tendency to over-estimate the CFE, hinting at the fact that ALIGNN may be capable of extrapolating across supercell sizes. These results suggest that the GNN models for CFE prediction could generalize across types of chemistries, structures, and system sizes for particular cases, but in general may need to be retrained and fine-tuned for specific datasets.
Active Learning (AL) Workflow
  1. A.
    Training the Ensemble of ALIGNN Models: We begin by training an ensemble of ALIGNN models to capture the variability and uncertainties associated with the predictions. To make this ensemble, we partition the original training set into multiple subsets, each containing a different combination of training, validation, and test data. A total of 100 different ALIGNN models have been trained, each on a unique subset of the data, allowing us to account for variability due to data partitioning. Figure S5(b) illustrates the ALIGNN predictions on the test dataset (we name it ALIGNN-1) vs. DFT calculations from 100 different ALIGNN models, highlighting the standard deviation in the predictions along with the mean.
  2. B.
    Prediction Across Expanded Defect Chemical Space: After training the ensemble of ALIGNN-1 models, we utilize them to predict the CFE of all the defects in the expanded chemical space. For each configuration, predictions are made using all 100 models in the ensemble, yielding a distribution of predictions. This approach enables us to not only obtain the mean prediction but also to quantify the uncertainty associated with each prediction. Figure S7(a) shows the violin plot of predicted mean CFE (averaged across 100 ALIGNN models) made across the entire set of defects.
  3. C.
    Uncertainty Quantification: The uncertainty of each prediction is quantified by analyzing the standard deviation of CFE among the predictions made by the 100 ALIGNN-1 models. In our AL framework, we employed the maximum uncertainty (MU) acquisition function Farache_Verduzco. The MU criterion is defined as $\text{MU}(x)=\sigma(x)$, where $\sigma(x)$ denotes the standard deviation (uncertainty) of the prediction. Figure S7(b) highlights the defects that maximize the MU acquisition function identified by the ALIGNN-1 models.
  4. D.
    Active Learning via Bayesian optimization and New DFT Calculations: We utilized Bayesian optimization to refine the predictions of the ALIGNN-1 models by selecting the 200 configurations that maximize the chosen acquisition function, prioritizing the most informative data points. These selected configurations were then used to launch new DFT calculations, ensuring that the model iteratively improves its accuracy and predictive performance. Although the model initially has not encountered certain defects, such as those in the CdxZn1-xTe composition, the predictions from ALIGNN-1 models are reasonable. Figure S7(c) shows ALIGNN-1 prediction (energy of the initial input defect structure, viz., unoptimized energy) vs. DFT calculation (unoptimized energy) for 200 selected defects based on acquisition function. To further improve the performance of ALIGNN-1 models, these new calculations are then incorporated into the training set, iteratively improving the accuracy of the model.
  5. E.
    Final Model Performance: Following the first iteration of the AL loop, we retrain ALIGNN-1 models and get new ALIGNN models (we name it ALIGNN-2) which are used to make predictions across the remaining defect space. The violin plot of predicted mean CFE (averaged across 100 ALIGNN-2 models) made on the remaining defect configurations is presented in Figure S7(d). The predictions are again evaluated using the acquisition function as shown in Figure S7(e), and 200 new configurations that maximize its value are selected for new DFT calculations. We observe a significant improvement in the ALIGNN predictions as depicted in Figure S7(f), closely matching the DFT-computed CFE across the selected defects. This rapid convergence after just one training cycle highlights the effectiveness of the AL approach in enhancing model performance, even in underexplored regions of the chemical space. Later on, newly obtained DFT data is again incorporated into the training set and we retrained the model (ALIGNN-3). Finally, the ALIGNN-3 models are used to predict the remaining unexplored defects.The rationale behind choosing the 200 defects that maximizes the MU acquisition function is driven by a careful consideration of our computational budget and capabilities.
Detailed Description of ALIGNN-Based Geometry Optimization
Figure  S8: Optimization of Cd vacancy (VCd) in CdTe using the trained ALIGNN model with two different optimization strategies: simulated annealing and Bayesian optimization.
Figure  S9: a) Simulated annealing optimization of a ClTe defect in CdSe0.50Te0.50 using ALIGNN-predicted crystal formation energy (CFE), showing energy minimization over successive steps. (b) Parity plot comparing ALIGNN-optimized CFE with DFT-optimized CFE across a set of defect structures, demonstrating strong agreement. (c) Parity plot comparing ALIGNN-unoptimized CFE with DFT-unoptimized CFE.
Figure  S10: Comparison of computational efficiency between a DFT-only workflow and the ALIGNN+SA+DFT approach for evaluating ClTe and AsTe defects in Cd0.50Zn0.50Te under Cd-rich conditions. Left: Defect formation energies (DFEs) as a function of Fermi level, computed using the HSE06+SOC functional on top of the PBE optimized structures. Right: Workflow comparison showing that direct DFT relaxation of all 108 symmetry-inequivalent AsTe configurations requires approximately 512 hours, while the ALIGNN+SA model identifies the lowest-energy site in minutes, followed by ALIGNN+PBE optimization. A single static HSE06+SOC calculation on the ALIGNN+PBE optimized configuration reduces the total computational time to just 8 hours.
Here, we combined the ALIGNN predictions with two commonly used optimization techniques to achieve energy minimization of new defect configurations:
  1. 1.
    Simulated Annealing (SA) Cheng_Gong, which is a gradient-free stochastic method employing random atomic perturbations guided by an annealing schedule.
  2. 2.
    Bayesian Optimization (BO) Cheng_Gong, which uses Gaussian processes to iteratively identify low energy configurations through atomic displacement exploration.
As an illustrative example, both simulated annealing and Bayesian optimization were applied to optimize a VCd defect within a $\times3$$\times$3 CdTe supercell (see Figure S8). Each method rapidly identified stable defect structures within minutes, significantly faster than traditional DFT methods. Simulated annealing notably discovered a Te–Te dimer configuration, which standard DFT often overlooks without prior chemical intuition.
Coupling ALIGNN prediction of CFE values with SA or BO enabled gradient-free energy minimization via atomic displacements applied within a cut-off radius around the defect center. These algorithms systematically searched for the lowest energy configuration using ALIGNN predictions at each step, allowing fast and guided optimization. As a case study, we applied both algorithms to optimize a Cd vacancy (VCd) defect in a $\times3$$\times$3 CdTe supercell, and the results are pictured in Figure  S8. Either method reaches a general energy convergence within 300 to 400 optimization steps which take a total of a few minutes to complete, but SA performs much better than BO: not only does it find a lower energy structure, but it actually discovers the configuration featuring a Te–Te dimer, which was reported by Kavanagh et al. sean_1 and which is easily missed by standard DFT optimization. To further evaluate the optimization capability of the ALIGNN model, we combined it with SA to optimize a variety of defects across the CdSexTe1-x and CdxZn1-xTe chemical space as shown in Figure  S9. For example, ClTe in Cd0.50Zn0.50Te was successfully optimized using ALIGNN+SA, as shown in Figure  S9 (a).
Table SVI: Performance of pretrained machine-learning force-field (MLFF) models applied directly to the PBE test dataset without retraining. The reported errors correspond to the root-mean-square error (RMSE) in crystal formation energy (CFE).
Model Pretrained Version RMSE (meV/atom)
MACE mace MACE_MPtrj_2022.9.model 93.79
M3GNet m3gnet M3GNet-MatPES-PBE-v2025.1-PES 62.10
CHGNet chgnet CHGNet-MatPES-PBE-2025.2.10-2.7M-PES 59.50
Figure S10 illustrates the DFEs of AsTe and ClTe in Cd0.50Zn0.50Te under Cd-rich conditions as a function of EF, computed using the HSE06+SOC functional on top of PBE-optimized lowest-energy sites. The right panel compares two workflows and highlights the computational advantage of incorporating ALIGNN+SA-based optimization. In the conventional DFT-only approach, all 108 symmetry-inequivalent AsTe configurations must be relaxed individually to identify the lowest-energy structure, requiring approximately 512 hours. In contrast, the ALIGNN+DFT workflow first uses the trained ALIGNN+SA model to rapidly evaluate and optimize the CFE for all configurations, identifying the most stable site in minutes. DFT calculations (PBE followed by HSE06+SOC) are then performed on the ALIGNN-predicted lowest-energy structure, reducing the total computational cost to just 8 hours. The next section presents an attempt to improve upon this by moving towards gradient-based optimization using force fields.
PBE Machine learning Force Field (MLFF) Model
M3GNet-based m3gnet MLFF models were trained using the energies, atomic forces, and stresses extracted from GGA–PBE calculation trajectories. Radial and three-body cutoffs were set to 6 Å and 6 Å, respectively. The loss was a weighted sum of RMSE for energies, forces, and stresses (weights = 1, 1, and 0.01, respectively). Training was performed on an NVIDIA A100 (80 GB) with batch size 64 and learning rate $5\times 10^{-4}$ until convergence. Geometry optimization with the MLFF used the FIRE algorithm in ASE ase_1, ase_2 with convergence criteria of mean atomic force $10^{-5}$ eV/Å or a maximum of 100 ionic steps. Models were trained separately for structures in five different charge states. To improve model accuracy on difficult configurations where predictions were poor, we used a two-stage training process:
  • Warm-up: In the first stage, we trained the model for a small number of epochs using uniform sampling. This helps the model develop a basic understanding of the data.
  • Error-aware reweighting: Next, we used the warm-up model to predict energies and forces for all training samples. Based on the prediction errors, we assigned a score to each sample. Samples with larger errors were considered harder. These error scores were then converted into sampling weights. Limits were applied to avoid extremely large or small weights and cap any outliers. These weights were passed to a WeightedRandomSampler in PyTorch, which increased the likelihood of selecting harder examples during training. The validation set remained unweighted. The reweighting step was repeated every 10–20 epochs to keep the weights up to date. This method helps the model focus more on difficult configurations while still learning broadly, which leads to better performance on complex regions of the data.
Figure S11(a–c) show parity plots for the M3GNet-MLFF models trained for charge states $q=+1$, $q=0$, and q=1q=-1. Models for the $q=+2$ and q=2q=-2 charge states are presented in Figure S12. Each parity plot compares DFT-computed CFE for test set points with the corresponding values from the MLFF prediction for different types of structures: bulk (pristine supercells without defects), defects (bulk supercells containing a single point defect or defect complex), defects at CdTe-ZnTe interfaces, and defects in CdTe dislocation core (DC). Overall, the MLFF predictions show remarkably low RMSE values for all types of bulk and defect configurations, similar to the ALIGNN models.
For $q=0$, the test prediction RMSE for bulk structures is 8.6 meV/atom, a similarly low value of 6.8 meV/atom for defect structures, and 3.7 meV/atom for interface. Dislocation core defect structures show slightly larger errors of >> 9.1 meV/atom, which is expected given their more complex local environments and atomic rearrangements. Similar errors were seen for charged defect structures as well. The overall agreement between DFT and MLFF predictions remains generally strong, indicating the robustness of the model across diverse chemical and structural environments.
One of the major advantages of having an MLFF model rather than a direct energy prediction model is the ability to use predicted atomic forces to perform geometry optimization based on gradient-based energy minimization, which is more computationally efficient than gradient-free optimization as shown in Figure S13. For example, we optimized the AsTe+ClTe defect in CdSe0.12Te0.88 using both ALIGNN (direct energy) and M3GNet (MLFF) and found that M3GNet achieved the optimization at a substantially lower computational cost compared to ALIGNN. Figure S11(d–f) show three different examples of using the MLFF model for optimizing challenging defect configurations: an AsTe substitutional defect in CdTe, a ZnCd defect in a CdTe-ZnTe interface structure, and an AsCd substitutional defect in a CdTe dislocation core structure. These cases represent chemically and structurally complex environments that are often found in devices with polycrystalline semiconductor thin films. The optimized q=+2,+1,0,1,2q=+2,+1,0,-1,-2 MLFF models were able to successfully capture local atomic rearrangements and produce low energy configurations consistent with DFT benchmarks. In each case, the MLFF-optimized configuration energy matches well with the DFT-optimized energy. Energy minimization is achieved in approximately 100 steps, with the entire relaxation process completing within a few minutes.
Figure  S11: Performance of M3GNet-MLFF models trained on the GGA-PBE dataset, shown in terms of predicted vs DFT crystal formation energy parity plots. The models were trained separately for (a) defect configurations with charge q=+1, (b) neutral q=0 defect and bulk configurations, and (c) defect configurations with charge q=-1. Here, "bulk" refers to pristine supercells without any defects, "defect" means bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "DC" indicates defects situated in CdTe dislocation core structures. Plots in (d–f) show the geometry optimization process taking into account 5 charge states for different example defect configurations: (d) a AsTe defect in CdTe, (e) a ZnCd defect at a CdTe/ZnTe interface, and (f) AsCd in a <111> CdTe dislocation core structure (DC).
Figure  S12: Parity plots for M3GNet-MLFF models trained on the GGA dataset, shown in terms of predicted vs actual (from DFT) crystal formation energies, trained separately for (a) defect configurations with charge q=+2, (b) defect configurations with charge q=-2. Here, "defect" represents bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "DC" indicates defects situated at CdTe dislocation core.
Figure  S13: (a) Optimization of an AsTe + ClTe complex in CdSe0.12Te0.88 using an M3GNET-MLFF model, and (b) comparison of the time taken by ALIGNN and the MLFF for optimizing selected defects in CdSe0.12Te0.88.
Figure  S14: Performance of the M3GNet-MLFF models trained on the augmented HSE06 dataset, shown in terms of predicted vs DFT crystal formation energy parity plots. The models were trained separately for (a) defect configurations with charge q=+1, (b) neutral q=0 defect and bulk configurations, and (c) defect configurations with charge q=-1. Here, "bulk" refers to pristine supercells without any defects, "defect" means bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "DC" indicates defects situated at CdTe dislocation core. Plots in (d–f) show the geometry optimization process taking into account 5 charge states for different example defect configurations: (d) an AsSe+ClSe defect complex in CdSe0.12Te0.88, (e) a Cdi defect at the CdTe/ZnTe interface, and (f) AsTe in <111> CdTe dislocation core structure.
Figure  S15: Parity plots for M3GNet-MLFF models trained on the HSE06 dataset, shown in terms of predicted vs actual (from DFT) crystal formation energies, trained separately for (a) defect configurations with charge q=+2, (b) defect configurations with charge q=-2. Here, "defect" represents bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "DC" indicates defects situated at CdTe dislocation core structure.
Figure  S16: Parity plot comparing DFT and MLFF-predicted crystal formation energies (CFE) for single and complex defects (q=0 charge state).
Figure  S17: Parity plots comparing MLFF-predicted defect formation energies with DFT values for different charge states and charge–correction schemes. Each panel shows the MLFF prediction plotted against the corresponding DFT value for charged defects with $q=+2$, +1+1, 1-1, and 2-2. The first row uses actual DFT charge corrections, the second row uses averaged charge corrections for each charge, and the third row applies no charge correction.
Figure  S18: Parity plot comparing MLFF-predicted and DFT-calculated charge-transition levels, with average charge correction value applied to MLFF prediction in each charge state. Four types of transition levels are calculated, namely +2/+1, +1/0, 0/-1, and -1/-2.
Figure  S19: Comparison of DFT- and MLFF-relaxed structures using PCA on SOAP descriptors dscribe_1, dscribe_2. Each point represents the structural fingerprint of a relaxed configuration projected onto a two-dimensional PCA space. Blue and orange markers correspond to DFT and MLFF optimizations, respectively. The near-overlapping distribution of DFT and MLFF points demonstrates that the latter accurately reproduces DFT-level structural relaxations across diverse bulk and defect configurations.
Figure  S20: The DeFecT-FF tool takes as input the bulk crystal, list of defects, bandgap, and VBM, performs MLFF-based geometry optimization, leverages an in-house chemical potential library to calculate charge-dependent energies $E(q)$ and defect formation energies $E_{f}$ as a function of the Fermi level $E_{F}$, and finally constructs $E_{f}$$E_{F}$ diagrams for defect thermodynamics. This tool is accessible via a nanoHUB web application.
Figure  S21: Illustration of the configuration search for AsSe defects. The left panel shows how a standard DFT relaxation can become trapped in a false ground state, while the right panel presents the relative energies of 14 symmetry-broken configurations obtained using the HSE-based MLFF. The true lowest-energy configuration (configuration 5) is highlighted in red.
Table SVII: Comparison of RMSE (eV) in MLFF defect formation energies for three correction schemes: (i) no charge correction, (ii) average charge-offset correction, and (iii) DFT charge correction applied to MLFF predictions.
Charge State No Correction Average Offset Correction DFT Charge Correction
q=2q=-2 0.30 0.23 0.20
q=1q=-1 0.29 0.21 0.19
$q=0$ 0.16 0.16 0.16
$q=+1$ 0.27 0.20 0.16
$q=+2$ 0.30 0.20 0.19
Charge Correction Using the Freysoldt Scheme
Charged defects in periodic boundary conditions introduce spurious electrostatic interactions between the charged defect, its periodic images, and the compensating background charge. To correct for these finite-size effects, we employed the Freysoldt Freysoldt2014-en correction scheme as implemented in the sxdefectalign code. The correction energy consists of two components: an image-charge term ($E_{\mathrm{PC}}$) and a potential alignment term ($q\,\Delta V$):
$E_{\mathrm{corr}}=E_{\mathrm{PC}}+q\,\Delta V.$ (4)
The image-charge correction is estimated using an isotropic model charge distribution screened by the static dielectric constant $\varepsilon$ of the host material:
$E_{\mathrm{PC}}=\frac{q^{2}\alpha}{2\,\varepsilon\,L},$ (5)
where $q$ is the defect charge state, $\alpha$ is the Madelung constant for the supercell geometry, and $L$ is the effective lattice parameter of the supercell. The potential alignment term $\Delta V$ is obtained from the planar-averaged electrostatic potential difference between the defective and bulk supercells in the far-field region:
$\Delta V=V_{\mathrm{defect}}^{\mathrm{far}}-V_{\mathrm{bulk}}.$ (6)